OpenStudy (anonymous):

Is 4x – 3 = 19 true, false, or open?

OpenStudy (anonymous):

4x-3=19 4x= 22 x= 22/4 = 6.5 It just means x=6.5 I have no idea if it is true, false or open. To me it simply do not make sense. A equation like: x-3= x-2 I can say it is false as (cancel the both x), 3 is not equals to 2. But for x=6.5 Because of the variable x, it is difficult to ascertain any truth value to it.

OpenStudy (anonymous):

open

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

2x =2x true

OpenStudy (anonymous):

what would 7 + 9 = –9 + 7 be then?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

16≠-2

OpenStudy (anonymous):

ohhh. okay thanks(:

OpenStudy (anonymous):

np